# Transcorrelated Theory with Pseudopotentials

### v1

@@ 1 Introduction
The transcorrelated (TC) method is an explicitly correlated method that has seen rapid development in recent years 7, 8, 9, 10, 11. TC takes a Jastrow factor, a function of interelectronic distances optimized in a first-quantized VMC calculation, and uses it to perform a similarity transformation on the second-quantized Hamiltonian. This transformation preserves the Hamiltonian’s eigenvalues while addressing the cusps and significantly improving basis set convergence 12, as well as compactifying the wave function 13. Although the transformation introduces challenging three-body terms, a recent approximation, xTC 14, removes the need to explicitly treat these terms, reducing the scaling in evaluation of the transcorrelated Hamiltonian by two orders of magnitude.
TC has shown very promising results in homogeneous electron gas (HEG) systems 15, 16, and the Hubbard model 13. It has also been applied to atoms and molecules with high accuracy using FCIQMC and CC methods 7, 17, 18, 19. TC has also been used to produce accurate results in simulations with quantum hardware, where the TC Hamiltonian enables the use of shallower circuit depths 20.
@@ 2.1 Pseudopotential approximation
Within the PP approximation the Coulombic interactions between the valence electrons and the atomic nuclei and core electrons are replaced with effective potentials. The PPs are constructed to reproduce the valence electron wave functions outside of a core region. This means that the electron-nucleus term of the Hamiltonian $\hat{H}_{en}$ is replaced by a sum over effective potentials $\hat{V}_{eff}$ for each atom:
$\hat{H}_{en}=\sum_{I}^{M}\sum_{i=1}^{N}\frac{Z_{I}}{|\mathbf{r}_{i}-\mathbf{R}_{I}|}\rightarrow\hat{H}^{PP}_{en}=\sum_{I}^{M}\sum_{j=1}^{N_{v}}\hat{V}^{I}_{eff}(|\mathbf{r}_{j}-\mathbf{R}_{I}|).$ (1)
@@
Above, $Z$ is the nuclear charge, $N$ ($M$) is the number of electrons (atoms), and $\mathbf{r}_{i}$ ($\mathbf{R}_{I}$) is the position of the $i$-th electron ( $I$th atom) with respect to the origin. With PPs, the summation runs over $N_{v}$ valence electrons.
@@
In the theory that follows, we focus on the effective potential of a single atom ($M=1$) for simplicity, but the extension to multiatom systems is straightforward.
@@
$V^{I}_{eff}$ should mimic the potential of the valence electrons due to the nucleus and the core electrons. It should also reproduce the exact electronic wave function outside of the core region specified by a cutoff radius $r_{c}$. $V_{eff}$ consists of a number of angular momentum channels, one local without spherical projections and one or many nonlocal channels, and is typically expressed as:
$\hat{V}_{eff}(r)=V_{lmax}(r)+\sum_{l=0}^{l_{max}-1}V_{l}(r)\sum_{m=-l}^{l}|Y_{lm}\rangle\langle Y_{lm}|.$ (2)
@@
Above, $Y_{lm}$ are the spherical harmonics, and $V_{l}(r)$ are the pseudopotential radial functions. $l_{max}$ is the maximum angular momentum quantum number included in the PP, which is also chosen as the local channel. $V_{l}$ are expressed as:
@@
$V_{l}(r)=\left\{\begin{array}[]{ll}-\frac{Z_{eff}}{r}\left(1-e^{-\alpha r^{2}}\right)+\alpha Z_{eff}re^{-\beta r^{2}}+\sum_{q=1}^{n}\gamma_{ql}e^{-\delta_{ql}r^{2}}&l=l_{max}\\ \sum_{q=1}^{m}\gamma_{ql}e^{-\delta_{ql}r^{2}}&l<l_{max}\end{array}\right.$ (3)
@@
We use two sets of PPs in this work: energy-consistent correlated electron PPs (eCEPPs) by Trail & Needs 21, and correlation-consistent effective core potentials (ccECPs) by Lubos-Mitas 22 et al. The eCEPPs have $l_{max}=2$ for the first-row atoms, with $n=4$ and $m=6$, while the ccECPs have $l_{max}=1$ and $n=m=1$, leaving only one gaussian term for the non-local channels. The different components of ccECPs and eCEPPs for C are shown in Fig. 1.
@@
Figure 1: $V_{l}(r)$ of eCEPPs (solid lines) and ccECPs (dashed lines) of C for different angular momentum components.
@@
The action of $\hat{V}_{eff}$ on a function $\phi(\mathbf{r})$ (with or without a Jastrow factor) of the position of electron $i$ is:
$\hat{V}_{eff}(r_{i})\phi(\mathbf{r}_{i})=\sum_{l=0}^{l_{max}}V_{l}(r_{i})\sum_{m=-l}^{l}Y_{lm}(\Omega_{\mathbf{r}_{i}})\int_{\left|\mathbf{r}{{}^{\prime}}\right|=r_{i}}d\Omega_{\mathbf{r}{{}^{\prime}}}\phi(\mathbf{r}{{}^{\prime}})Y_{lm}(\Omega_{\mathbf{r}{{}^{\prime}}}).$ (4)
@@
Because $Y_{lm}(\phi=0,\theta=0)=0$ for $m\neq 0$, we can simplify this expression – by choosing the $z$-axis to be along $\mathbf{r}_{i}$ – to 23:
$\hat{V}_{eff}(r_{i})\phi(\mathbf{r}_{i})=\sum_{l=0}^{l_{max}}V_{l}(r_{i})Y_{l0}(\Omega_{\mathbf{r}_{i}})\int_{\left|\mathbf{r}{{}^{\prime}}\right|=r_{i}}d\Omega_{\mathbf{r}{{}^{\prime}}}\phi(\mathbf{r}{{}^{\prime}})Y_{l0}(\Omega_{\mathbf{r}{{}^{\prime}}}),$ (5)
@@ 2.2 Pseudopotentials in the Transcorrelated Hamiltonian
In transcorrelation theory the Hamiltonian is supplemented with a Jastrow factor $J$, which is a function to describe interparticle correlations. The Jastrow factor in this work is expressed as:
$\begin{split}J(\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{N})&=\sum_{i=1}^{N}\chi(\mathbf{r}_{i})+\sum_{i<j}u(\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|)+\sum_{i<j}f(\mathbf{r}_{i},\mathbf{r}_{j},\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|)\\ &=\sum_{i>j}J_{2}(\mathbf{r}_{i},\mathbf{r}_{j})\end{split}$ (6)
including $1$-body (electron-nucleus), $2$-body (electron-electron), and $3$-body (electron-electron-nucleus) terms. We have simplified the summations by including the sum over one-body terms in the two-body sum 14.
@@
$\begin{split}[\hat{O},J]=&\sum_{k=1}^{N}\sum_{i<j}\left[\hat{O}(\mathbf{r}_{k}),J_{2}(\mathbf{r}_{i},\mathbf{r}_{j})\right]=\sum_{k=1}^{N}\sum_{i<j}\Pi_{ij}^{k}(\hat{O}),\\ [[\hat{O},J],J]=&\sum_{k=1}^{N}\sum_{i<j}\sum_{l<m}\left[\left[\hat{O}(\mathbf{r}_{k}),J_{2}(\mathbf{r}_{i},\mathbf{r}_{j})\right],J_{2}(\mathbf{r}_{l},\mathbf{r}_{m})\right]=\sum_{k=1}^{N}\sum_{i<j}\sum_{l<m}\Gamma^{k}_{ijlm}(\hat{O}).\end{split}$ (8)
@@
With PPs, the $V_{eff}$ terms may not commute with the Jastrow factor. In addition, commutator summation in the similarity transformation is not guaranteed to terminate at the second commutator. To estimate the effect of the PP on the transcorrelated Hamiltonian, we make an approximation of only considering the $2$-body terms of the PP commutators, ignoring the sum over $i>j>m$ in Eq. (9). This approximation is made under the assumption that the $3$-body interactions close to the core are neglible. Explicit treatment of the $3$-body terms, possibly under the xTC approximation, is left for future work in case it is needed. Within this approximation, the transcorrelated second-quantized Hamiltonian becomes
@@
The terms $h$ and $V$ are traditional one- and two-body terms of the second-quantized Hamiltonian, while the calculation of the transcorrelated components $K$ and $L$ arising from the kinetic energy operator has been described elsewhere 7, 14. The pseudopotential terms $P$ are computed as:
$\begin{split}P^{pq}_{rs}=\Bigg{\langle}\phi_{p}\phi_{q}\left|\Pi^{1}_{12}(\hat{H}^{PP}_{en})+\Pi^{2}_{12}(\hat{H}^{PP}_{en})+\frac{1}{2}\Gamma^{1}_{1212}(\hat{H}^{PP}_{en})+\frac{1}{2}\Gamma^{2}_{1212}(\hat{H}^{PP}_{en})\right|\phi_{r}\phi_{s}\Bigg{\rangle}\end{split}$ (11)
@@
$\begin{split}\Pi^{i}_{ij}(\hat{H}^{PP}_{en})=&\hat{H}^{PP}_{en}(\mathbf{r}_{i})J_{2}(\mathbf{r}_{i},\mathbf{r}_{j})-J_{2}(\mathbf{r}_{i},\mathbf{r}_{j})\hat{H}^{PP}_{en}(\mathbf{r}_{i}),\\ \Gamma^{i}_{ijij}(\hat{H}^{PP}_{en})=&\hat{H}^{PP}_{en}(\mathbf{r}_{i})J_{2}(\mathbf{r}_{i},\mathbf{r}_{j})^{2}-2J_{2}(\mathbf{r}_{i},\mathbf{r}_{j})\hat{H}^{PP}_{en}(\mathbf{r}_{i})J_{2}(\mathbf{r}_{i},\mathbf{r}_{j})\\ +&J_{2}(\mathbf{r}_{i},\mathbf{r}_{j})^{2}\hat{H}^{PP}_{en}(\mathbf{r}_{i}).\end{split}$ (12)
Therefore, in order to evaluate the PP commutators (under the present approximation of restricting ccECP corrections to two-body terms) in Eq. (11), we need to apply Eq. (5) to calculate terms of the type $\hat{H}^{PP}_{en}\phi$, $\hat{H}^{PP}_{en}J_{2}\phi$, and $\hat{H}^{PP}_{en}J_{2}^{2}\phi$. This allows us to construct the transcorrelated Hamiltonian in Eq. (10). We have included also higher-order terms in the PP commutators in Eq. (12) in our calculations, under the assumption that the $3$-body terms are negligible.
@@ 3 Calculations
In this work we study the use of pseudopotentials with the transcorrelated Hamiltonian in atoms Be, B, C, N, O, and F, as well as their $+1$ ions. We also study the total and atomization energies of molecules CN, CO, CF, N2, O2, F2, H2O and CO2. We use the aug-cc-pVD/T/QZ basis sets (AVXZ, with X=D,T,Q), optimized individually for eachs pseudopotential21, 22. Ioniszation energies with ccECPs are evaluated with non-augmented cc-pVD/T/QZ basis sets (PVXZ, with X=D,T,Q). .
The geometries of the molecules are taken from those in the HEAT database 24. The Hartree-Fock (HF) calculations are done with the PYSCF code 25. The Jastrow factors used are of Drummond-Towler-Needs type 26 and are optimised with respect to the variance of the Hartree-Fock wave function using the VMC method of the CASINO package 27.

### v2

@@ 1 Introduction
The transcorrelated (TC) method is an explicitly correlated method that has seen rapid development in recent years 7, 8, 9, 10, 11, 12. TC takes a Jastrow factor, a function of interelectronic distances optimized in a first-quantized VMC calculation, and uses it to perform a similarity transformation on the second-quantized Hamiltonian. This transformation preserves the Hamiltonian’s eigenvalues while addressing the cusps and significantly improving basis set convergence 13, as well as compactifying the wave function 14, 12. Although the transformation introduces challenging three-body terms, a recent approximation, xTC 15, removes the need to explicitly treat these terms, reducing the scaling in evaluation of the transcorrelated Hamiltonian by two orders of magnitude.
TC has shown very promising results in homogeneous electron gas (HEG) systems 16, 17, and the Hubbard model 14. It has also been applied to atoms and molecules with high accuracy using FCIQMC and CC methods 7, 18, 19, 20. TC has also been used to produce accurate results in simulations with quantum hardware, where the TC Hamiltonian enables the use of shallower circuit depths 21.
@@ 2.1 Pseudopotential approximation
Within the PP approximation the Coulombic interactions between the valence electrons and the atomic nuclei and core electrons are replaced with effective potentials. The PPs are constructed to reproduce the valence electron wave functions outside of a core region. This means that the electron-nucleus term of the Hamiltonian $\hat{H}_{\text{en}}$ is replaced by a sum over effective potentials $\hat{V}_{\text{eff}}$ for each atom:
$\hat{H}_{\text{en}}=\sum_{I}^{M}\sum_{i=1}^{N}\frac{Z_{I}}{|\mathbf{r}_{i}-\mathbf{R}_{I}|}\rightarrow\hat{H}^{\text{PP}}_{\text{en}}=\sum_{I}^{M}\sum_{j=1}^{N_{v}}\hat{V}^{I}_{\text{eff}}(|\mathbf{r}_{j}-\mathbf{R}_{I}|).$ (1)
@@
Above, $Z$ is the nuclear charge, $N$ ($M$) is the number of electrons (atoms), and $\mathbf{r}_{i}$ ($\mathbf{R}_{I}$) is the position of the $i$-th electron ($I$-th atom) with respect to the origin. With PPs, the summation runs over $N_{v}$ valence electrons.
@@
In the theory that follows, we focus on the effective potential of a single atom ($M=1$) for to simplify presentation, but the extension to multiatom systems is straightforward.
@@
Outside of the core region $V^{I}_{\text{eff}}$ should mimic the potential felt by the valence electrons due to the nucleus and the core electrons. It should also reproduce the exact electronic wave function outside of the core region specified by a cutoff radius $r_{c}$. $V_{\text{eff}}$ consists of a number of angular momentum channels, one local without spherical projections and one or more nonlocal channels:
$\hat{V}_{\text{eff}}(r)=V_{l_{\text{max}}}(r)+\sum_{l=0}^{l_{\text{max}}-1}V_{l}(r)\sum_{m=-l}^{l}|Y_{lm}\rangle\langle Y_{lm}|.$ (2)
@@
Above, $Y_{lm}$ are the spherical harmonics, and $V_{l}(r)$ are the pseudopotential radial functions. $l_{\text{max}}$ is the maximum angular momentum quantum number included in the PP, which is also chosen as the local channel. $V_{l}$ are expressed as:
@@
$V_{l}(r)=\left\{\begin{array}[]{ll}-\frac{Z_{\text{eff}}}{r}\left(1-e^{-\alpha r^{2}}\right)+\alpha Z_{\text{eff}}re^{-\beta r^{2}}+\sum_{q=1}^{n}\gamma_{ql}e^{-\delta_{ql}r^{2}}&l=l_{\text{max}}\\ \sum_{q=1}^{m}\gamma_{ql}e^{-\delta_{ql}r^{2}}&l<l_{\text{max}}\end{array}\right.$ (3)
@@
We use two sets of PPs in this work: energy-consistent correlated electron PPs (eCEPPs) by Trail & Needs 22, and correlation-consistent effective core potentials (ccECPs) by Lubos-Mitas 23 et al. The eCEPPs have $l_{\text{max}}=2$ for the first-row atoms, with $n=4$ and $m=6$, while the ccECPs have $l_{\text{max}}=1$ and $n=m=1$, leaving only one gaussian term for the non-local channels. The different components of ccECPs and eCEPPs for C, N, O, and F are shown in Fig. 1. In the figure, it can be seen that the eCEPPs have much smaller potential absolute values than the ccECPs.
@@
Figure 1: $V_{l}(r)$ of eCEPPs (solid lines) and ccECPs (dashed lines) of C, N, O, and F. Black dashed lines show the Coulombic potential of the nucleus.
@@
The action of $\hat{V}_{\text{eff}}$ on a function $f(\mathbf{r})$ (with or without a Jastrow factor) of the position of electron $i$ is:
$\hat{V}_{\text{eff}}(r_{i})f(\mathbf{r}_{i})=\sum_{l=0}^{l_{\text{max}}}V_{l}(r_{i})\sum_{m=-l}^{l}Y_{lm}(\Omega_{\mathbf{r}_{i}})\int_{\left|\mathbf{r}{{}^{\prime}}\right|=r_{i}}d\Omega_{\mathbf{r}{{}^{\prime}}}f(\mathbf{r}{{}^{\prime}})Y_{lm}(\Omega_{\mathbf{r}{{}^{\prime}}}).$ (4)
@@
Because $Y_{lm}(f=0,\theta=0)=0$ for $m\neq 0$, we can simplify this expression – by choosing the $z$-axis to be along $\mathbf{r}_{i}$ – to 24:
$\hat{V}_{eff}(r_{i})f(\mathbf{r}_{i})=\sum_{l=0}^{l_{max}}V_{l}(r_{i})Y_{l0}(\Omega_{\mathbf{r}_{i}})\int_{\left|\mathbf{r}{{}^{\prime}}\right|=r_{i}}d\Omega_{\mathbf{r}{{}^{\prime}}}f(\mathbf{r}{{}^{\prime}})Y_{l0}(\Omega_{\mathbf{r}{{}^{\prime}}}),$ (5)
@@ 2.2 Pseudopotentials in the Transcorrelated Hamiltonian
In transcorrelation theory, the Hamiltonian is supplemented with a Jastrow factor $J$, which is a function to describe interparticle correlations. With $N$ electrons and $M$ nuclei in a system, the Jastrow factor in this work is of Drummond-Towler-Needs (DNT) type 25:
$\begin{split}J(\{\mathbf{r}_{i}\},\{\mathbf{R}_{I}\})&=\sum_{I=1}^{M}\sum_{i=1}^{N}\chi(\mathbf{r}_{iI})+\sum_{i<j}u(\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|)+\sum_{I=1}^{M}\sum_{i<j}f(\mathbf{r}_{iI},\mathbf{r}_{jI},\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|)\\ &=\sum_{I=1}^{M}\sum_{i>j}J_{2}(\mathbf{r}_{i},\mathbf{r}_{j};\mathbf{R}_{I}),\end{split}$ (6)
including $1$-body ($\chi,$electron-nucleus), $2$-body ($u$, electron-electron), and $3$-body ($f$, electron-electron-nucleus) terms. Each of the terms is optimized for interparticle distances under a chosen cutoff value, characteristic of the DNT Jastrow form. Above, $\mathbf{r}_{iI}=\mathbf{r}_{i}-\mathbf{R}_{I}$. In the above, in order to simplify the derviation of the pseudopotential commutator equations 15, we have included all of the terms involving $\chi,u$ and $f$ into the term $J_{2}(\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{R}_{I})$. This approach gives TC contributions to the $2$- and $3$-body terms $\langle pr|qs\rangle$ and $\langle pqr|stu\rangle$ of the Hamiltonian, including also the $\chi$-term contribution to these terms. We call this approach the ”combined” Jastrow treatment. Alternatively, one could treat the $1$- and $2$-body terms in the Jastrow factor separately (which we call the ”separate” Jastrow treatment) and include the contribution of $\chi$ into the $1$-body terms $\langle p|q\rangle$. The separate Jastrow treatment does not remove the contribution of the $\chi$-term from the $2$-body integrals, however, because of the presence of a cross-term (see Supplementary Material). With small $\chi$ cutoff values the combined and separated treatments yield almost exactly the same coupled cluster total energies for atoms and molecules at equilibrium geometries, with discrepancy of <.$<.1$mHa. However, we found difficulties with the separated Jastrow treatment with PPs whenever the Jastrow cutoff values for the $\chi$ functions exceeded the inter-nuclear bond lengths, and for this reason we chose the combined approach in this study. We present the equations for the separated treatment of the $1$- and $2$-body terms in the Jastrow factor in the Supplementary Material.
@@
$\begin{split}[\hat{O},J]=&\sum_{k=1}^{N}\sum_{i<j}\left[\hat{O}(\mathbf{r}_{k}),\sum_{I}^{M}J_{2}(\mathbf{r}_{i},\mathbf{r}_{j};\mathbf{R_{I}})\right]=\sum_{k=1}^{N}\sum_{i<j}\Pi_{ij}^{k}(\hat{O}),\\ [[\hat{O},J],J]=&\sum_{k=1}^{N}\sum_{i<j}\sum_{l<m}\left[\left[\hat{O}(\mathbf{r}_{k}),\sum_{I}^{M}J_{2}(\mathbf{r}_{i},\mathbf{r}_{j};\mathbf{R_{I}})\right],\sum_{J}^{M}J_{2}(\mathbf{r}_{l},\mathbf{r}_{m};\mathbf{R_{J}})\right]=\sum_{k=1}^{N}\sum_{i<j}\sum_{l<m}\Gamma^{k}_{ijlm}(\hat{O}).\end{split}$ (8)
@@
With PPs, the $V_{\text{eff}}$ terms do not commute with the Jastrow factor. In addition, commutator summation in the similarity transformation is not guaranteed to terminate at the second commutator. To estimate the effect of the PP on the transcorrelated Hamiltonian, we make an approximation of only considering the $2$-body terms of the PP commutators, ignoring the sum over $i>j>m$ in Eq. (9). This approximation is made under the assumption that the $3$-body interactions of the valence electrons close to the core are negligible. Explicit treatment of the $3$-body terms, possibly under the xTC approximation, is left for future work in case it is needed. Within this approximation, the transcorrelated second-quantized Hamiltonian becomes
@@
The terms $h$ and $V$ are traditional one- and two-body terms of the second-quantized Hamiltonian, while the calculation of the transcorrelated components $K$ and $L$ arising from the kinetic energy operator has been described elsewhere 7, 15. The pseudopotential terms $P$ are computed as:
$\begin{split}P^{pq}_{rs}=\Bigg{\langle}\phi_{p}\phi_{q}\left|\Pi^{1}_{12}(\hat{H}^{\text{PP}}_{\text{en}})+\Pi^{2}_{12}(\hat{H}^{\text{PP}}_{\text{en}})+\frac{1}{2}\Gamma^{1}_{1212}(\hat{H}^{\text{PP}}_{\text{en}})+\frac{1}{2}\Gamma^{2}_{1212}(\hat{H}^{\text{PP}}_{\text{en}})\right|\phi_{r}\phi_{s}\Bigg{\rangle}\end{split}$ (11)
@@
$\begin{split}\Pi^{i}_{ij}(\hat{H}^{\text{PP}}_{\text{en}})=&\sum_{I}^{M}\left[\hat{H}^{\text{PP}}_{\text{en}}(\mathbf{r}_{i})J_{2}(\mathbf{r}_{i},\mathbf{r}_{j};\mathbf{R_{I}})-J_{2}(\mathbf{r}_{i},\mathbf{r}_{j};\mathbf{R_{I}})\hat{H}^{\text{PP}}_{\text{en}}(\mathbf{r}_{i})\right],\\ \Gamma^{i}_{ijij}(\hat{H}^{\text{PP}}_{\text{en}})=&\hat{H}^{\text{PP}}_{\text{en}}(\mathbf{r}_{i})\left[\sum_{I}^{M}J_{2}(\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{R_{I}})\right]^{2}-2\sum_{I}^{M}J_{2}(\mathbf{r}_{i},\mathbf{r}_{j};\mathbf{R_{I}})\hat{H}^{\text{PP}}_{\text{en}}(\mathbf{r}_{i})\sum_{J}^{M}J_{2}(\mathbf{r}_{i},\mathbf{r}_{j};\mathbf{R_{J}})\\ +&\left[\sum_{I}^{M}J_{2}(\mathbf{r}_{i},\mathbf{r}_{j};\mathbf{R_{I}})\right]^{2}\hat{H}^{\text{PP}}_{\text{en}}(\mathbf{r}_{i}).\end{split}$ (12)
Therefore, in order to evaluate the PP commutators (under the present approximation of restricting ccECP corrections to two-body terms) in Eq. (11), we need to apply Eq. (5) to calculate terms of the type $\hat{H}^{\text{PP}}_{\text{en}}\phi$, $\hat{H}^{\text{PP}}_{\text{en}}J_{2}\phi$, and $\hat{H}^{\text{PP}}_{\text{en}}J_{2}^{2}\phi$. This allows us to construct the transcorrelated Hamiltonian in Eq. (10). We have included also higher-order terms in the PP commutators in Eq. (12) in our calculations, under the assumption that the $3$-body terms are negligible.
@@ 3 Calculations
3.1 Evaluation of transcorrelated Hamiltonian with pseudopotentials